OpenStudy (anonymous):

What is the percent decrease from 5600 to 3160?

OpenStudy (anonymous):

43.57

hero (hero):

@Kashan, were you planning on providing steps with that? Providing the answer alone doesn't help anyone.

OpenStudy (anonymous):

@Hero sure..

hero (hero):

Percent change formula: a = 1st number b = 2nd number \[\%\Delta = \frac{|a-b|}{a} \times 100\]

hero (hero):

@Kashan, I don't need the steps, @mitchelsewbaran needs them.

OpenStudy (anonymous):

first find the percentage of 3160 from 5600 by. (3160/5600) x 100 this will give the increasing percentage... subtract that percentage from 100 and it will give you the decreasing percentage..

OpenStudy (anonymous):

is it 21%?

OpenStudy (anonymous):

No... first solve (3160/5600) x 100 what you got..??

OpenStudy (anonymous):

44%

OpenStudy (anonymous):

(3160/5600) x 100 = 0.5643 x 100 = 56.43 % now subtract it from 100 you'll get your desired answer

OpenStudy (anonymous):

43.57